[Brahms-l] Shift report 20040320 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Mar 20 2004 - 07:59:08 EST
------------------------------------------------------------
Date: Sat Mar 20 07:59:08 2004, Shift: 00:00-08:00
Supervisor: Michael Murray, with: Chaitanya Kalavagunta
Log book#15 pages 92 - 92

Shift Summary:
------------------------------------------------------------
1) Very long store. There was a problem with AGS magnet and MCR cannot refill RHIC.
2) Three data runs and one short run for ZDC triggers
3) T5-1 lost 2 plains
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 20 07:59:49 2004

This archive was generated by hypermail 2.1.8 : Sat Mar 20 2004 - 08:00:07 EST